OpenStudy (anonymous):

School guidelines require there must be at least 2 chaperones per 25 students going on a school trip. How many chaperones must there be for 55 students?

OpenStudy (anonymous):

Well, you can't have a part of a chaperon, so how many do you think?

OpenStudy (anonymous):

I think 5- because 25 times 2 is 50, it will be at least 4.

OpenStudy (anonymous):

Yah. Because half a block of students makes sense . And if you look at it as a fraction, you want something \(\gt \dfrac{2}{25}=.08 \) Well, \(\dfrac{5}{55} = \dfrac{1}{11} \approx .09\) so more than .08.

OpenStudy (anonymous):

And that should be \(\ge \dfrac{2}{25}\)